### v2

# Practical Policy Distillation for Reinforcement Learning in Radio Access Networks

Abstract
[…]
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This is the author’s version of the paper accepted for publication in IEEE International Symposium on Personal, Indoor and Mobile Radio Communications, 2025. The final published version is available at IEEE Xplore, https://doi.org/10.1109/PIMRC62392.2025.11274591.
ACK
acknowledgement
AI
artificial intelligence
ARQ
automatic repeat request
BE
baseline encoding
BC
behavioral cloning
BCQ
batch-constrained deep Q-learning
CCTR
Channel-Conditioned Target Return
CNN
convolutional neural network
CQL
conservative Q-learning
CT
continuous time
CV
computer vision
DAVG
discounted average
DLLA
downlink link adaptation
DP
dynamic programming
DQN
deep Q-network
DT
decision transformer
eMBB
enhanced mobile broadband
FB
full buffer
ILLA
inner-loop link adaptation
gNB
next generation NodeB
gNB-CU
gNB centralized unit
gNB-DU
gNB distributed unit
GPI
generalized policy iteration
LA
link adaptation
LSTM
long short-term memory
LT
learnable time
KL
Kullback-Leibler
MBB
mobile broadband
MC
Monte Carlo
MDP
Markov decision process
ML
machine learning
MLP
multi-layer perceptron
mMIMO
massive MIMO
mMTC
massive machine-type communications
NACK
negative acknowledgement
NLP
natural language processing
OLLA
outer-loop link adaptation
PE
positional encoding
PHY
physical layer
RL
reinforcement learning
RLC
radio link control
RNN
recurrent neural network
RTGs
returns-to-go
RTG
return-to-go
RvS
reinforcement learning via supervised learning
SACo
state-action coverage
SADCo
state-action density coverage
SCSU
single-cell single-user
SE
spectral efficiency
TBS
transport block size
TD
temporal difference
TQ
relative trajectory quality
URLLC
ultra-reliable low-latency communications
VAE
variational auto-encoder
2G
Second Generation
3G
[…]
3GPP
[…]
4G
[…]
5G
[…]
AA
Antenna Array
AC
Admission Control
AD
Attack-Decay
ADSL
Asymmetric Digital Subscriber Line
AHW
Alternate Hop-and-Wait
AMC
Adaptive Modulation and Coding
AoA
angle of arrival
AP
Access Point
APA
Adaptive Power Allocation
AR
autoregressive
ARMA
Autoregressive Moving Average
ATES
Adaptive Throughput-based Efficiency-Satisfaction Trade-Off
AWGN
additive white Gaussian noise
$B$
Branch and Bound
BD
Block Diagonalization
BER
bit error rate
BF
Best Fit
BLER
block error rate
BPC
Binary power control
BPSK
Binary Phase-Shift Keying
BPA
Best pilot-to-data power ratio (PDPR) Algorithm
BRA
Balanced Random Allocation
BCRB
Bayesian Cramér-Rao Bound
BS
base station
CAP
Combinatorial Allocation Problem
CAPEX
Capital Expenditure
CBF
Coordinated Beamforming
CBR
Constant Bit Rate
CBS
Class Based Scheduling
$C$
Congestion Control
CDF
cumulative distribution function
CDMA
Code-Division Multiple Access
CL
Closed Loop
CLPC
Closed Loop Power Control
CNR
Channel-to-Noise Ratio
CPA
Cellular Protection Algorithm
CPICH
Common Pilot Channel
CoMP
Coordinated Multi-Point
CQI
channel quality indicator
CRB
Cramér-Rao Bound
CRM
Constrained Rate Maximization
CRN
Cognitive Radio Network
CS
Coordinated Scheduling
CSI
channel state information
CSIR
channel state information at the receiver
CSIT
channel state information at the transmitter
CUE
cellular user equipment
D2D
device-to-device
DCA
Dynamic Channel Allocation
DE
Differential Evolution
DFT
Discrete Fourier Transform
DIST
Distance
DL
downlink
DMA
Double Moving Average
DMRS
demodulation reference signal
D2DM
D2D Mode
DMS
D2D Mode Selection
DPC
Dirty Paper Coding
DRA
Dynamic Resource Assignment
DSA
Dynamic Spectrum Access
DSM
Delay-based Satisfaction Maximization
ECC
Electronic Communications Committee
EFLC
Error Feedback Based Load Control
EI
Efficiency Indicator
eNB
Evolved Node B
EPA
Equal Power Allocation
EPC
Evolved Packet Core
EPS
Evolved Packet System
ESPRIT
estimation of signal parameters via rotational invariance
E-UTRAN
Evolved Universal Terrestrial Radio Access Network
ES
Exhaustive Search
FDD
frequency division duplexing
FDM
Frequency Division Multiplexing
FER
Frame Erasure Rate
$F$
Fast Fading
FIM
Fisher information matrix
FSB
Fixed Switched Beamforming
FST
Fixed SNR Target
FTP
File Transfer Protocol
GA
Genetic Algorithm
GBR
Guaranteed Bit Rate
GLR
Gain to Leakage Ratio
GOS
Generated Orthogonal Sequence
GPL
GNU General Public License
GRP
Grouping
HARQ
hybrid automatic repeat request
HMS
Harmonic Mode Selection
HOL
Head Of Line
HSDPA
High-Speed Downlink Packet Access
HSPA
High Speed Packet Access
HTTP
HyperText Transfer Protocol
ICMP
Internet Control Message Protocol
ICI
Intercell Interference
ID
Identification
ISAC
integrated sensing and communication
IEEE
Institute of Electrical and Electronics Engineers
IETF
Internet Engineering Task Force
ILP
Integer Linear Program
JRAPAP
Joint RB Assignment and Power Allocation Problem
UID
Unique Identification
IID
Independent and Identically Distributed
IIR
Infinite Impulse Response
ILP
Integer Linear Problem
IMT
International Mobile Telecommunications
INV
Inverted Norm-based Grouping
IoT
Internet of Things
IP
Internet Protocol
IPv6
Internet Protocol Version 6
ISD
Inter-Site Distance
ISI
Inter Symbol Interference
ITU
International Telecommunication Union
JOAS
Joint Opportunistic Assignment and Scheduling
JOS
Joint Opportunistic Scheduling
JP
Joint Processing
JS
Jump-Stay
KKT
Karush-Kuhn-Tucker
L3
Layer-3
LAC
Link Admission Control
LC
Load Control
LOS
Line of Sight
LP
Linear Programming
LS
least squares
LTE
Long Term Evolution
LTE-A
LTE-Advanced
LTE-Advanced
Long Term Evolution Advanced
M2M
Machine-to-Machine
MAC
Medium Access Control
MANET
Mobile Ad hoc Network
MCS
modulation and coding scheme
MDB
Measured Delay Based
MDI
Minimum D2D Interference
MF
Matched Filter
MG
Maximum Gain
MH
Multi-Hop
MIMO
multiple input multiple output
MINLP
Mixed Integer Nonlinear Programming
MIP
Mixed Integer Programming
MISO
Multiple Input Single Output
MLE
maximum likelihood estimator
MLWDF
Modified Largest Weighted Delay First
MME
Mobility Management Entity
MMSE
minimum mean squared error
MOS
Mean Opinion Score
MPF
Multicarrier Proportional Fair
MRA
Maximum Rate Allocation
MR
Maximum Rate
MRC
Maximum Ratio Combining
MRT
Maximum Ratio Transmission
MRUS
Maximum Rate with User Satisfaction
MS
mobile station
MSE
mean squared error
MSI
Multi-Stream Interference
MTC
Machine-Type Communication
MTSI
Multimedia Telephony Services over IMS
MTSM
Modified Throughput-based Satisfaction Maximization
MU-MIMO
multiuser multiple input multiple output
MU
multi-user
MUSIC
multiple signal classification
NAS
Non-Access Stratum
NB
Node B
NE
Nash equilibrium
NCL
Neighbor Cell List
NLOS
Non-Line of Sight
NMSE
Normalized Mean Square Error
NORM
Normalized Projection-based Grouping
NP
Non-Polynomial Time
NR
New Radio
NRT
Non-Real Time
NSPS
National Security and Public Safety Services
O2I
Outdoor to Indoor
OFDMA
orthogonal frequency division multiple access
OFDM
orthogonal frequency division multiplexing
OFPC
Open Loop with Fractional Path Loss Compensation
O2I
Outdoor-to-Indoor
OL
Open Loop
OLPC
Open-Loop Power Control
OL-PC
Open-Loop Power Control
OPEX
Operational Expenditure
ORB
Orthogonal Random Beamforming
JO-PF
Joint Opportunistic Proportional Fair
OSI
Open Systems Interconnection
PAIR
D2D Pair Gain-based Grouping
PAPR
Peak-to-Average Power Ratio
P2P
Peer-to-Peer
PC
Power Control
PCI
Physical Cell ID
PDF
probability density function
PDPR
pilot-to-data power ratio
PER
Packet Error Rate
PF
Proportional Fair
P-GW
Packet Data Network Gateway
PL
Pathloss
PPR
pilot power ratio
PRB
physical resource block
PROJ
Projection-based Grouping
ProSe
Proximity Services
PS
Packet Scheduling
PSAM
pilot symbol assisted modulation
PSO
Particle Swarm Optimization
PZF
Projected Zero-Forcing
QAM
Quadrature Amplitude Modulation
QoS
Quality of Service
QPSK
Quadri-Phase Shift Keying
RAISES
Reallocation-based Assignment for Improved Spectral Efficiency and Satisfaction
RAN
radio access network
RAT
Radio Access Technology
RATE
Rate-based
RB
resource block
RBG
Resource block broup
REF
Reference Grouping
RM
Rate Maximization
RNC
Radio Network Controller
RND
Random Grouping
RRA
Radio Resource Allocation
RRM
radio resource management
RSCP
Received Signal Code Power
RSRP
Reference Signal Receive Power
RSRQ
Reference Signal Receive Quality
$R$
Round Robin
RRC
Radio Resource Control
RSSI
Received Signal Strength Indicator
RT
Real Time
RU
Resource Unit
RUNE
RUdimentary Network Emulator
RV
Random Variable
SAC
Session Admission Control
SCM
Spatial Channel Model
SC-FDMA
Single Carrier - Frequency Division Multiple Access
SD
Soft Dropping
S-D
Source-Destination
SDPC
Soft Dropping Power Control
SDMA
Space-Division Multiple Access
SER
Symbol Error Rate
SES
Simple Exponential Smoothing
S-GW
Serving Gateway
SINR
signal-to-interference-plus-noise ratio
SI
Satisfaction Indicator
SIP
Session Initiation Protocol
SISO
single input single output
SIMO
Single Input Multiple Output
SIR
signal-to-interference ratio
SLNR
Signal-to-Leakage-plus-Noise Ratio
SMA
Simple Moving Average
SNR
signal-to-noise ratio
SORA
Satisfaction Oriented Resource Allocation
SORA-NRT
Satisfaction-Oriented Resource Allocation for Non-Real Time Services
SORA-RT
Satisfaction-Oriented Resource Allocation for Real Time Services
SPF
Single-Carrier Proportional Fair
SRA
Sequential Removal Algorithm
SRS
Sounding Reference Signal
SSB
synchronisation signal block
SU-MIMO
single-user multiple input multiple output
SU
Single-User
SVD
Singular Value Decomposition
TCP
Transmission Control Protocol
TDD
time division duplexing
TDMA
Time Division Multiple Access
TETRA
Terrestrial Trunked Radio
TP
Transmit Power
TPC
Transmit Power Control
TTI
transmission time interval
TTR
Time-To-Rendezvous
TSM
Throughput-based Satisfaction Maximization
TU
Typical Urban
UE
user equipment
UEPS
Urgency and Efficiency-based Packet Scheduling
UL
uplink
UMTS
Universal Mobile Telecommunications System
URI
Uniform Resource Identifier
URM
Unconstrained Rate Maximization
UT
user terminal
VR
Virtual Resource
VoIP
Voice over IP
WAN
Wireless Access Network
WCDMA
Wideband Code Division Multiple Access
WF
Water-filling
WiMAX
Worldwide Interoperability for Microwave Access
WINNER
Wireless World Initiative New Radio
WLAN
Wireless Local Area Network
WMPF
Weighted Multicarrier Proportional Fair
WPF
Weighted Proportional Fair
WSN
Wireless Sensor Network
WWW
World Wide Web
XIXO
(Single or Multiple) Input (Single or Multiple) Output
ZF
zero-forcing
ZMCSCG
Zero Mean Circularly Symmetric Complex Gaussian
I Introduction
[…]
AI is increasingly recognized as a key enabler of improved performance and efficiency in RANs. AI-driven solutions have been explored across a broad range of RAN functions, including channel estimation and symbol decoding at the physical layer [12, 17]; fast control loops, such as link adaptation and radio resource management, at the medium access layer [11, 6, 5, 15]; and slower functionalities such as mobility management, load balancing, and network configuration [2, 9]. Factors such as user mobility, interference, traffic fluctuations, and the heterogeneity of both RAN infrastructure and user equipment (UE) introduce significant variability. As a result, achieving reliable performance in real-world RAN deployments requires AI models that can adapt to diverse and dynamic network conditions [16]. In other words, models must generalize effectively and perform robustly, even on previously unseen data [19].
Training an AI model for a RAN function to generalize across diverse RAN environments—rather than fine-tuning separate models for specific conditions, such as those encountered in a single radio cell—is essential for scalable AI integration in RANs [16]. Reducing model proliferation enables more efficient and cost-effective lifecycle management. In contrast, models fine-tuned to specific network conditions often lack the robustness required to adapt to changes in network and environmental dynamics, necessitating frequent updates and compromising performance consistency. However, achieving effective model generalization requires extensive and diverse training data, as well as sufficiently large models capable of learning and extrapolating complex patterns from large datasets [19].
Although RANs are data-rich systems, a fundamental challenge in deploying AI within existing infrastructure is the computational constraint of legacy baseband hardware. Designed primarily for efficient signal processing and protocol execution, 4G and 5G baseband units were not dimensioned to support large AI models. As a result, a mismatch often arises between the demands of AI-based network optimization and the limited hardware resources available for inference in RAN systems. This restricts the size and complexity of models that can be practically deployed, thereby limiting their ability to generalize effectively.
Knowledge distillation [8] has been proposed to address the challenge of deploying complex models on resource-constrained devices. It enables knowledge transfer from a large, high-performing teacher model to a smaller, computationally efficient student model, with negligible degradation in accuracy [10]. Originally introduced for supervised learning tasks [4, 10], recent work [13, 18] has extended these techniques to policy distillation in reinforcement learning for control applications. However, the ability of policy distillation to retain the generalization capability of the teacher model—and its applicability to practical problems in RANs—remains relatively unexplored.
This paper explores policy distillation as a method to replace traditional RAN functions with AI-based alternatives under realistic RAN operational constraints. Our goal is to distill AI models that comply with RAN baseband hardware constraints while retaining the sophisticated decision-making capabilities of large, computationally intensive teacher models. Focusing on the link adaptation (LA) function, we employ domain randomization to generate diverse training datasets from varied simulated environments. These datasets are then used to train compact, computationally efficient student models that closely replicate the behavior of high-capacity teacher models. We evaluate two distillation strategies: single-policy distillation, which compresses a single scenario-agnostic teacher into a compact student model; and multi-policy distillation, which integrates knowledge from multiple scenario-specific teachers into a generalist student model. Experimental evaluations in a 5G-compliant simulator confirm both strategies dramatically reduce model size while preserving strong generalization.
II The Hidden Challenges of AI in RAN
Significant research has been devoted to AI-driven solutions for various RAN functions, examining algorithms, models, training methods, and data collection strategies. However, limited attention has been paid to the real-time processing constraints that hinder the deployment of AI in practical RAN settings. Existing RAN baseband hardware—typically based on FPGA, ASIC, or DSP architectures optimized for specialized, repetitive signal processing tasks (e.g., FFT, modulation/demodulation) and protocol execution—was not originally designed to support AI workloads, such as tensor computations or floating-point operations. Consequently, current hardware lacks the flexibility and computational headroom (e.g., memory and processing resources) required to effectively support large AI models. Additionally, baseband units commonly operate at full computational capacity under strict latency and power constraints, necessitating minimal—ideally zero—additional overhead from any AI integration.
Another crucial constraint is the stringent latency requirement of critical user-plane functions in Layer 1 (L1) and Layer 2 (L2), such as resource scheduling and link adaptation. These functions must execute within tens of microseconds—for example, link adaptation (LA) typically requires less than ten microseconds—thereby significantly limiting permissible inference latency. These latency constraints restrict AI model complexity, memory footprint, and per-inference computational cost, often limiting feasible AI models to sizes no larger than a few tens of kilobytes due to limited on-chip memory and strict inference-time budgets. Collectively, these constraints underscore the need for efficient AI deployment strategies tailored specifically to legacy RAN infrastructure.
To address these challenges, we investigate policy distillation—a technique that transfers knowledge from large, computationally intensive AI models to smaller models suitable for resource-constrained hardware. We apply this approach to LA, an L2-critical-loop RAN function executed multiple times (once per scheduled UE) on a sub-millisecond timescale.
III Problem formulation
We consider the problem of LA in the downlink of a RAN, where the transmission rate is dynamically adjusted to match the link capacity by optimizing the modulation and coding scheme (MCS) parameters based on time- and frequency-varying channel conditions and radio interference. In 5G systems, the MCS parameters—comprising the modulation order and code rate—are represented by an MCS index [1], denoted by integer values $m$ ranging from 0 to M1M-1.
LA algorithms deployed in present-day RANs typically rely on two control loops: an inner loop that maps the latest channel state information (CSI) reported by a UE to an initial signal-to-interference-plus-noise ratio (SINR) estimate, and an outer loop that adjusts this estimate based on hybrid automatic repeat request (HARQ) feedback to maximize spectral efficiency while meeting a target block error rate (BLER). However, the outer loop converges slowly [3] to a fixed BLER target; therefore, it struggles to adapt to dynamic channel conditions under bursty traffic. In contrast, reinforcement learning (RL)-based approach can learn a policy that predicts the optimal MCS index in real time, offering the potential to overcome the limitations of rule-based LA algorithms.
III-A RL-based Link Adaptation
We model the LA problem as an episodic Markov decision process (MDP) $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right\rangle$, where each episode corresponds to the lifespan of a UE data packet—from its first transmission to either successful reception or packet drop after $N$ transmission attempts, as illustrated in Fig. 1. The action space $\mathcal{A}=\{a\mid a\in\{0,\dots,M-1\}\}$ comprises $M$ discrete values, representing the MCS indices used to adapt the link spectral efficiency. In our evaluations, we consider 28 MCS indices (i.e., $M=28$), as defined in Table 5.1.3.1-2 of the 5G specifications [1].
Figure 1: Example of MDP episode definition for LA.
Following [7], we use a reward function $r_{n}=r(s_{n},a_{n})\in\mathcal{R}$ for each individual packet transmission, defined as
[…]
where $n\in\{0,\dots,N-1\}$ is the packet transmission index, $\alpha\in\mathbb{R}_{+}$ is a penalty coefficient for retransmissions, and $SE_{n}$ denotes the spectral efficiency achieved upon successful reception in the $n^{\mathrm{th}}$ transmission attempt.
To support model generalization across the RAN environment during training, we consider two key enablers. First, we design the state $s\in\mathcal{S}$ to encapsulate a mixture of semi-static and dynamic information. The semi-static information serves two purposes: (i) to characterize the heterogeneity of the RAN deployment and its configuration relative to the UE—including deployment type, location, orientation, interrelations among network sites or radio cells, and parameters such as antenna array type, carrier frequency, bandwidth, and transmit power—and (ii) to describe the UE’s characteristics (e.g., chipset and capabilities), enabling the model to operate across a diverse set of devices. The dynamic information, in contrast, captures state variables relevant to LA behavior, such as path loss, CSI, HARQ feedback, and many more; see, e.g., [7] for the complete list of states.
Additionally, we leverage a distributed training architecture combined with domain randomization to improve model robustness by exposing the agent to varied and diverse environmental conditions. Domain randomization introduces controlled stochasticity into the training environment to bridge the simulation-to-reality gap, as further described in Section V.
IV Policy Distillation
To address the challenge of achieving model generalization while adhering to the constraints imposed by RAN hardware, we merge knowledge distillation with domain randomization.
Knowledge distillation, referred to as policy distillation in the RL context [13], involves transferring a policy learned by an agent (the teacher), through interaction with an environment, into a smaller, simpler model (the student) without significant performance degradation. Policy distillation can be performed in both online and offline settings [13, 14, 18]. In online policy distillation, the student model learns concurrently with the teacher, as the latter improves its policy through ongoing interaction with the environment. In contrast, offline policy distillation assumes that the teacher has been fully trained. The student is then trained using a dataset generated by the teacher’s fixed policy.
In this paper, we only focus on the offline approach, assuming a trained deep Q-network (DQN) policy for LA as the teacher, following the design presented in Section III.
IV-A Offline Policy Distillation
Following [14], we consider a distillation dataset $\mathcal{D}^{T}=\{(s_{i},\mathbf{q}_{i}^{T})\}_{i}$, where each sample $i$ consists of a state $s_{i}\in\mathcal{S}$ from the LA function and a vector of unnormalized Q-values $\mathbf{q}_{i}^{T}=\pi^{T}(s_{i})$ produced by the teacher policy $\pi^{T}$. These Q-values serve as target outputs for training a student policy $\pi^{S}$.
To guide the training, we employ the Kullback-Leibler (KL) divergence loss, defined as
$L_{\mathrm{KL}}(\mathcal{D}^{T},\pi^{S})=\sum_{i=1}^{\mid\mathcal{D}^{T}\mid}\mathrm{softmax}\left(\frac{\mathbf{q}^{T}_{i}}{\tau}\right)\ln\frac{\mathrm{softmax}(\frac{\mathbf{q}^{T}_{i}}{\tau})}{\mathrm{softmax}(\mathbf{q}^{S}_{i})}\;.$ (2)
Here, a softmax transformation is applied to both the teacher’s Q-values $\mathbf{q}_{i}^{T}$, scaled by a temperature parameter $\tau>0$, and the student’s Q-values $\mathbf{q}_{i}^{S}$, which are left unscaled. In the context of policy distillation, Rusu et al. [14] recommend using lower values of $\tau$ to sharpen the output distribution. Since Q-values represent the expected future discounted rewards for each possible action, sharpening accentuates the differences in action preferences and improves the fidelity of the transferred policy.
Unlike prior work, our objective is to distill a student policy $\pi^{S}$ that retains the teacher’s generalization capability while being significantly more compact. To this end, we revisit and extend two offline policy distillation strategies: single- and multi-policy distillation [14].
IV-B Generalization via Single-Policy Distillation
Our first approach combines single-policy distillation with domain randomization. In this case, we distill a student policy $\pi^{S}$ from the knowledge of a single expert-level teacher $\pi^{T}$. However, to retain the teacher’s ability to generalize across the RAN environment, the distillation dataset $\mathcal{D}^{T}$ must capture the teacher’s behavior under diverse radio communication conditions. To this end, we also apply domain randomization when generating the distillation samples $(s_{i},\mathbf{q}_{i}^{T})$, as described in Section V, ensuring that the state samples $s_{i}$ are drawn from heterogeneous network deployments with varying configurations, radio environments, traffic and interference types, load conditions, user device types, and so on.
Practical RAN deployments are, by nature, a source of extreme diversity due to the variety of radio environments, network configurations, traffic patterns, and interference distributions encountered across radio cells. Therefore, if a teacher model is trained using the collective experience of distributed actors operating in different radio cells, a diverse distillation dataset $\mathcal{D}^{T}$ can be constructed by reusing data stored in the teacher’s replay memory. Specifically, the trained teacher can be evaluated on state observations $s_{i}$ already present in the replay memory used during training.
IV-C Generalization via Multi-Policy Distillation
Our second approach is inspired by multi-policy distillation. Training a large RL teacher model from the collective experience of radio cells across a live RAN can be costly and may negatively impact RAN performance due to the exploration process. As an alternative path to achieving model generalization under hardware constraints, we propose training multiple teacher models using less invasive methods and applying multi-policy distillation to induce generalization by fusing their policies into a single student model.
Specifically, we consider $N$ independently trained teacher policies $\pi^{T_{j}}$, each specialized—rather than generalized—for a specific network environment or deployment. This allows each teacher model to be dimensioned according to the hardware constraints of the target RAN. For instance, different teacher models may be trained separately using drive tests in specific network contexts (e.g., urban, rural, high-speed), as well as with proprietary networks, lab testbeds, or high-fidelity network simulators.
After training a set of teacher models, we generate the corresponding distillation datasets, denoted as $\mathcal{D}^{T_{j}}=\{(s_{i},\mathbf{q}_{i}^{T_{j}})\}_{i=0}^{D_{j}}$ for all $j\in\{0,\dots,N-1\}$, following the same methodology as in single-policy distillation. We then shuffle and aggregate these datasets to train a unified student model. Since the student learns from the combined behavior of multiple specialized teachers, the resulting policy generalizes effectively across diverse RAN environments.
V Numerical examples
To evaluate the potential of knowledge distillation for facilitating the deployment of AI in RANs, we focus on its ability to reduce model size to meet the hardware limitations of current RAN technology while achieving the performance and retaining the generalization capability of a larger teacher model. To this end, we consider both single-policy and multi-policy distillation applied to a RL policy for MCS index selection in downlink LA, and compare the performance of a larger teacher model against multiple student models of varying size across a range of performance metrics. To assess whether the student models retain the generalization ability of the teacher, we evaluate them using three representative benchmark scenarios not encountered during training. Furthermore, we demonstrate that training a small model directly with RL—rather than applying policy distillation—fails to achieve comparable performance and generalization.
V-A Training Setup
[…]
Parameter Value range
Antenna array $1\times 2\times 2$ MIMO (4)
$8\times 4\times 2$ mMIMO (64)
[…]
Number of sub-bands {51, 106, 133, 217, 273}
[…]
UE antennas {2, 4}
Maximum TX rank {2, 4}
Maximum DL TX 5
Number FB UEs {1, 5, 10}
Number MBB UEs {5, 10, 25, 50, 100, 150}
[…]
Indoor probability {0.2, 0.4, 0.8}
TABLE I: Parameters for domain randomization.
[…]
V-B Testing Setup
To assess model generalization, we evaluate all models against three benchmark scenarios not explicitly encountered during training: single-cell single-user (SCSU), multiple input multiple output (MIMO), and massive MIMO (mMIMO), each assuming three cells with full buffer (FB) traffic and 10 UEs.
We begin by evaluating the teacher model against the LA baseline commonly adopted in present-day RANs, as described in Section IV. In our benchmark scenarios, the teacher achieves an average UE throughput gain of $12\%$ and a spectral efficiency (SE) improvement of $8\%$ over the baseline, demonstrating its ability to perform well in diverse scenarios not seen during training. However, since our primary objective is to assess whether policy distillation can retain the teacher’s generalization capability while significantly reducing model size, we focus our analysis on comparing student performance with that of the corresponding teacher across the three benchmark scenarios.
We consider three performance metrics: average UE throughput ($T$), BLER, and episodic reward ($r$). The results are reported using the relative gain/loss. Values close to zero indicate that the student models closely approximate the teacher’s behavior. For example, the relative throughput gain/loss ($\Delta T$) is computed as:
$\Delta T=100\times\frac{T_{\mathrm{student}}-T_{\mathrm{teacher}}}{T_{\mathrm{teacher}}}\;.$
Analogous definitions apply to $\Delta\mathrm{\acs{BLER}}$ and $\Delta r$. Finally, to assess whether the distillation process effectively transfers the teacher’s policy, we compare the probability density function (PDF) of the actions returned by the teacher and the corresponding student models in each benchmark scenario.
Distillation Student Model MIMO Scenario mMIMO Scenario SCSU Scenario
$\Delta T$ $\Delta\mathrm{BLER}$ $\Delta r$ $\Delta T$ $\Delta\mathrm{BLER}$ $\Delta r$ $\Delta T$ $\Delta\mathrm{BLER}$ $\Delta r$
Single-policy Student $4\times 64$ $0.04\% +$$4.7\%$ $5\% 0.0$$02\%$ $2.4\%$ $0.8\%$ $0.21\%$ $2.1\%$ $0.8\%$
Student $4\times 32$ $0.14\% +$$2.3\%$ $0.9\%$ $15\% 0.0$$0\%$ $1.2\%$ $0.02\%$ $6.6\%$ $1.1\%$
Student $3\times 32$ $0.17\%$ $6.5\%$ $1.7\%$ $0.09\% +$$4.7\%$ $1.7\%$ $0.54\%$ $6.5\%$ $2.2\%$
Multi-policy Student $4\times 64$ $1.2\%$ $4.8\%$ $2.5\%$ $1.0\%$ $2.7\%$ $0.8\%$ $1.6\%$ $2.5\%$ $0.8\%$
Student $4\times 32$ $2.3\%$ $2.9\%$ $3.3\%$ $1.9\%$ $3.7\%$ $1.2\%$ $1.9\%$ $2.5\%$ $1.6\%$
Student $3\times 32$ $2.8\%$ $5.7\%$ $3.7\%$ $2.3\%$ $4.0\%$ $1.5\%$ $2.2\%$ $0.0\%$ $1.9\%$
TABLE II: Relative gain/loss (%) of student models with respect to their teacher models, across performance metrics and benchmark scenarios.
V-C Generalization via Single-Policy Distillation
We apply single-policy distillation to an RL policy for MCS index selection in downlink LA, trained to generalize across the RAN environment. To this end, we combine the previously introduced distributed RL training architecture with domain randomization to improve model resilience to uncertainties in the RAN environment, such as network deployment and configuration, traffic conditions, and user population.
Specifically, we train the teacher model using 5000 simulations, distributed across 16 actors, each with randomized network parameters based on Table I. Each simulation lasts 3 seconds and consists of one to three tri-sectorial radio sites, randomly configured as MIMO or mMIMO, according to antenna characteristics in Table I. Site parameters such as inter-site distance, cell radius, bandwidth, and downlink transmit power are also randomly sampled from Table I. Further randomization includes cell loads, traffic types, indoor/outdoor UE ratios, and UE receiver types. UEs are generated with random counts of FB and mobile broadband (MBB) traffic, with traffic patterns modeled from real-world data. UE configurations—including the number of antenna elements, speed, and receiver type—are randomized to reflect hardware and algorithmic differences across devices (e.g., CSI estimation).
To enable the distillation process to retain the generalization capability of the teacher model, we further apply domain randomization to generate the distillation dataset $\mathcal{D}^{T}$ by testing the teacher against 1000 randomized network simulations, resulting in approximately 20 million samples—providing sufficient data diversity for effective distillation.
(a) MIMO
(b) mMIMO
(c) SCSU
Figure 2: Comparison of the average throughput achieved by each student relative to the teacher in the MIMO, mMIMO and SCSU benchmark scenarios.
Figure 2 and Table II (single-policy rows) demonstrate that distilling a teacher model (7 layers of 128 units) into progressively smaller student models—from 4 layers of 64 units down to 3 layers of 32 units, representing approximately a 30-fold compression—can effectively preserve performance. Specifically, throughput degradation remains within $1\%$ across the MIMO, mMIMO, and SCSU scenarios; BLER deviation is limited to within $\pm 7\%$ relative to the teacher model, with the mid-sized student model (4 layers of 32 units) even achieving a $6.6\%$ improvement in BLER for SCSU. Moreover, the episodic reward declines by no more than $1.7\%$. In contrast, the control model (3 layers of 32 units), trained from scratch, significantly underperforms—by approximately $12\%$ in MIMO, $10\%$ in mMIMO, and $25\%$ in SCSU—as indicated by the purple markers falling clearly below the 0% reference line in Fig. 2.
Figure 3 further confirms that the distilled student models effectively replicate the teacher model’s behavior, both in terms of throughput and the actions selected. Specifically, Fig. 3(a)–(c) illustrate a near-perfect overlap between the teacher (blue) and student models (orange, green, red) in the cumulative distribution function (CDF) of UE throughput. In contrast, the smaller control model trained from scratch (purple) deviates considerably—especially in the SCSU scenario. Similarly, Fig. 3(d)–(f) confirm that the student models closely match the teacher’s distribution of MCS indices. The control model, however, collapses onto a limited subset of indices, displaying pronounced peaks near indices 10 and 16, and rarely selecting indices above 19.
Collectively, these results confirm that single-policy distillation, when combined with domain randomization, effectively preserves the teacher model’s generalization capability, even after substantial compression—approximately 30-fold. Directly training a similarly sized model, however, leads to poor generalization performance.
V-D Generalization via Multi-Policy Distillation
Table II (multi-policy rows) shows that a single student model, distilled from the combined knowledge of three scenario-specific teacher models, closely retains the performance of the original experts despite an approximate 30-fold parameter reduction (to a model with 3 layers of 32 units). Across the three benchmark environments, the throughput deficit of the smallest student model never exceeds $2.8\%$ (MIMO scenario), while the larger student models maintain throughput within $1.0\%$ to $2.3\%$. Changes in BLER remain modest, ranging from an improvement of $4.0\%$ (mMIMO scenario) to a maximum increase of $5.7\%$ (MIMO scenario), never exceeding a +6% deviation. The maximum observed reduction in episodic reward is $3.7\%$. Collectively, these tight performance bounds indicate that multi-policy distillation effectively compresses and integrates specialized knowledge from multiple scenario-specific expert models into a single lightweight generalist model capable of robustly generalizing across diverse benchmark conditions.
(a) MIMO
(b) mMIMO
(c) SCSU
(d) MIMO
(e) mMIMO
(f) SCSU
Figure 3: Comparison of the teacher and students policies across the three benchmark scenarios in terms of: CDF of the UE throughput (a)-(c); and actions distribution, represented by the PDF of selected MCS values (d)-(f).
VI Conclusions
In this paper, we demonstrated that policy distillation is a viable approach to replacing traditional RAN functions with AI-based counterparts that conform to the computational and memory constraints of existing baseband hardware. Focusing on the LA function, we applied domain randomization to assemble richly varied training sets from multiple simulated RAN scenarios. These datasets enabled us to train lightweight student models that faithfully imitate the decision logic of high-capacity teacher models. We assessed two distillation paradigms: (i) single-policy distillation, which compresses a unified, scenario-agnostic teacher into a compact student; and (ii) multi-policy distillation, which merges the expertise of several scenario-specific teachers into a single generalist student. Through experiments, we showed that both strategies substantially reduce model size—bringing their footprints well within existing hardware limits—while retaining near-teacher performance and broad generalization across diverse network conditions. Our findings underscore policy distillation as a practical, minimally invasive pathway for embedding AI-driven RAN functions into current deployments.